# The Power of LLM-Generated Synthetic Data for Stance Detection in Online Political Discussions

### v1

@@ Abstract
Stance detection holds great potential for enhancing the quality of online political discussions, as it has shown to be useful for summarizing discussions, detecting misinformation, and evaluating opinion distributions. Usually, transformer-based models are used directly for stance detection, which require large amounts of data. However, the broad range of debate questions in online political discussion creates a variety of possible scenarios that the model is faced with and thus makes data acquisition for model training difficult. In this work, we show how to leverage LLM-generated synthetic data to train and improve stance detection agents for online political discussions: (i) We generate synthetic data for specific debate questions by prompting a Mistral-7B model and show that fine-tuning with the generated synthetic data can substantially improve the performance of stance detection. (ii) We examine the impact of combining synthetic data with the most informative samples from an unlabelled dataset. First, we use the synthetic data to select the most informative samples, second, we combine both these samples and the synthetic data for fine-tuning. This approach reduces labelling effort and consistently surpasses the performance of the baseline model that is trained with fully labeled data. Overall, we show in comprehensive experiments that LLM-generated data greatly improves stance detection performance for online political discussions.
@@ 1 Introduction
With the recent advent of powerful generative Large Language Models (LLMs) such as ChatGPT [^1] https://chat.openai.com/, Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023), investigating how synthetic data from LLMs can be leveraged for downstream tasks has become a very important and relevant topic of research. It is well known that Natural Language Processing (NLP) tasks based on single-label or multi-label classification such as stance and sentiment detection benefit immensely from large amounts of labelled data (Mehrafarin et al., 2022). However, in many cases, labelled data is scarce, especially in the realm of online political discussions: one issue is that gathering labelled data may not be allowed for privacy reasons. Another issue is that especially in stance detection for online political discussions comments and their stances are intrinsically linked to a question. There are a potentially infinite number of specific questions open for debate for which different types of comments (and their stances) arise. This can be a problem, since the samples may not be varied or pertain to only one class, hence usually yielding an unbalanced dataset. For instance, performance drops severely for questions in the test dataset, that are not present in the training dataset (Vamvas and Sennrich, 2020).
An obvious approach to stance detection might be to apply LLMs directly, which has been tried with mixed results (see Section 2). Furthermore, LLMs are not yet fast enough at inference time for smaller real world interactive deployment in online discussions, unless adequate hardware infrastructure is provided. Also, using public hosting infrastructure may not be feasible when use cases require anonymity and privacy.
@@
Figure 1: We investigate the use of LLM-generated synthetic data for stance detection in online political discussions. (A) We generate synthetic data  for specific questions using a Mistral-7B model. The synthetic data is then used to fine-tune the stance detection model. We show that fine-tuning with synthetic data improves the performance of the model, since the synthetic data is roughly faithful to the real data’s  underlying distribution. However, some real world samples  cannot be captured by the synthetic data. (B) We therefore use the synthetic data to identify the most informative samples  in the unlabelled real data pool. That is, we identify the samples that are on the boundary of the two classes of the synthetic data which are better off labelled by human experts. We show that combining the synthetic data with the manually labelled most informative samples improves the performance of the model even further.
While LLMs might not be directly useful for stance detection, we show in this paper, that they can be used to generate synthetic training data for more specific (and smaller) NLP models, such as BERT (Devlin et al., 2019), to create powerful stance detection for online political discussions that potentially work for any type of question. How exactly synthetic data from an LLM should be combined with a small pool of unlabelled training data will be investigated along the following questions:
Question 1 (Q1): Does fine-tuning with synthetic data improve the stance detection model? Answer: Yes, and the more the better!
One problem of stance detection datasets for online political discussions is the lack of (balanced) data for specific questions (sometimes also called issues). While fine-tuning with synthetic data is possible for BERT-based models, it was previously found to be challenging (Mahmoudi et al., 2024). Using more powerful LLMs, we can show that fine-tuning with synthetic data does improve stance detection performance in online political discussions. This also holds when augmenting existing labelled data.
Question 2 (Q2): Can the labelling effort be reduced by controlling what example to manually label using active learning with synthetic data? Answer: Yes, even if we don’t have real true labels initially.
Another challenge in stance detection for online political discussions is labelling data, which is expensive and time consuming. Active learning can help us to identify the most informative samples that are then labelled manually, even if there are initially no labelled samples. For this we use the synthetic samples which have labels by construction, to select the most ambiguous (and thus most informative) samples in the unlabelled dataset.
Question (Q3): Can fine-tuning using the most informative samples plus the synthetic data outperform fine-tuning with a completely manually labelled dataset? Answer: Yes, it can!
Our study reveals that the most effective way to improve stance detection performance is to combine both synthetic data and the most informative samples obtained via active learning. Surprisingly, this is even better than using a full dataset that has been completely manually labelled.
@@ 2 Background
Stance detection is a sub-task of sentiment analysis (Romberg and Escher, 2023) and opinion mining (ALDayel and Magdy, 2021) that aims to automatically detect the stance (e.g., favor, against, or neutral) of the author of a contribution towards a discussed issue, or target. In online political discussions, stances can be defined as the side that the author is on, e.g., for or against an increase in value added tax. Research has identified stance detection as an important yet challenging task; it is important because the automated detection of stances can, e.g., improve discussion summarisation (Chowanda et al., 2017), facilitate the detection of misinformation (Hardalov et al., 2022), and provide more comprehensive evaluations of opinion distributions in online political discussions and participation processes (Romberg and Escher, 2023). Stance detection is also implemented as a feature in many recommender systems and discussion platforms (Küçük and Can, 2020). Still, due to its dependency on context, stance detection is a highly challenging task. In online political discussions, for example, detecting the stance of a contribution towards a question requires a proper identification of both the question under discussion and the position of the contributor towards it. However, users regularly deviate from the original question and discuss multiple topics in the same thread (Ziegele et al., 2014), resulting in a high number of potential questions with often very few instances for training purposes. Fine-tuning transformer-based models (Vaswani et al., 2017) to solve stance detection is a common practice, but training these models requires a large amount of annotated data, which for the large variety of questions in online political discussions is unfeasible to acquire. We therefore in our work, show how to improve stance detection for online political discussions with synthetic data.
Active Learning.
@@
The aim of active learning is to minimize the effort of labelling data, while simultaneously maximizing the model’s performance. This is achieved by selecting a query strategy that chooses the most interesting samples from a set of unlabelled data points, which we refer to as most informative samples. These samples are then passed to, e.g., a human annotator for labelling. The usual setup consists of iteratively training a model on a small subset of manually labelled data that stems from a pool of unlabelled data. There exist many different query strategies such as Query By Comittee (QBC, (Seung et al., 1992)), Minimum Expected Entropy (MEE, Holub et al. (2008) or Contrastive Active Learning (CAL, Margatina et al. (2021)). By actively choosing samples and asking for the correct labelling, the model is able to learn from few labelled data points, which is advantageous especially when annotated datasets are not available. Within the domain of political text analysis, many different tasks lack large amounts of annotated data. It has been already shown in the past that these tasks can benefit from the active learning: e.g., stance detection (Kucher et al., 2017), topic modeling (Romberg and Escher, 2022), speech act classification (Schmidt et al., 2023) or toxic comment classification (Miller et al., 2020). In this work, we examine how LLM-generated synthetic data can be used instead of real labelled data to select the most informative samples to be manually labelled.
Using LLM-generated synthetic data for fine tuning.
@@
Recent work has shown that synthetic data generated from LLMs can be used to improve the performance of a model on downstream tasks. Møller et al. (2023) showed that synthetic data can be used to improve the performance of a model on downstream classification tasks by comparing the performance of a model finetuned on LLM-generated data to crowd annotated data. In many cases the model finetuned on LLM-generated data outperforms the model finetuned on crowd annotated data. Mahmoudi et al. (2024) study the use of synthetic data for data augmentation in stance-detection. The authors use GPT-3 to generate synthetic data for a specific topic with mixed results due to the inability of GPT-3 to generate good data samples. In our work,we use a newer LLM model, Mistral-7B, which generates better synthetic data samples and show that we can generate synthetic data that matches the real data distribution. Veselovsky et al. (2023) analyse in which ways synthetic data is best generated for tasks like sarcasm detection and sentiment analysis. The authors reach the conclusion that grounding the prompts to generate the synthetic data to real samples helps improve the quality of the synthetic data. Similarly, Li et al. (2023) argue that subjectivity in the classification task determines whether synthetic data can be used effectively.
Using LLMs directly for stance detection.
It has been shown that LLMs can be used directly for stance detection such as (Cruickshank and Xian Ng, 2023), (Burnham, 2023). However, the general conclusion of these studies is that while LLMs are competitive with other transformer models such as BERT, especially for edge cases, they exhibit replication issues. (Burnham, 2023) also discuss the posibility of pre-training models on more specific data to improve the generalisation capability of the model. We therefore focus on using LLMs to generate synthetic data to solve key challenges in stance detection such as the lack of available data for specific topics and labelling large amounts of data, rather than using LLMs directly for the task.
@@ 3 Method
In line with our declared contributions, we present our core ideas to improve the performance of the stance detection model: (i) To fine-tune the model with synthetic data, we first define the synthetic dataset and show our pipeline to generate it. The baseline model is then further fine-tuned on the synthetic data. (ii) We then present our synthetic data-based approach to identify the most informative samples from a batch of unlabelled data. We propose a synthetic extension to the QBC (Query by Commitee, (Seung et al., 1992)) method, where the synthetic data act as an ensemble of experts.
@@ 3.1 Preliminaries
Political discussions are typically centered around questions $q\in\mathcal{Q}$ (sometimes also called issues or targets). For stance detection, we usually have for each of these questions $q$ a set of labelled data $\mathcal{D}^{(q)}=\{(x^{(i)},y^{(i)})\}_{i=1}^{I}$ where $x^{(i)}\in\mathcal{X}$ is a statement (or comment) and $y^{(i)}$ is the stance of the statement, with $y^{(i)}\in\{0,1\}=\mathcal{Y}$. Note, that we use the notation $\mathcal{D}^{(q)}$ for labelled and for unlabelled datasets (then the labels are ignored). We view the stance detection model as a binary classification function $f:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{Y}$, where we included the question as input to provide context. The classifier $f$ can be split into two parts, $f=h\circ g$: the first part is a pre-trained language model $g:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{E}$ such as BERT (Devlin et al., 2019) or RoBERTa (Liu et al., 2019). The second part is a classifier $h:\mathcal{E}\rightarrow\mathcal{Y}$ on top of the pre-trained model $g$. The pre-trained language model $g$ outputs an embedding $e^{(i)}\in\mathcal{E}$ given a question $q$ and an input $x^{(i)}$. The classifier $h$ is usually a linear layer or a multi-layer perceptron, which takes the embedding $e^{(i)}$ from $g$ as input and outputs a prediction $\hat{y}^{(i)}$. The stance detection model is fine-tuned by minimizing the cross-entropy loss between the predicted labels $\hat{y}^{(i)}=f(q,x^{(i)})$ and the actual labels $y^{(i)}$.
@@ 3.2 Generating Synthetic Data for Stance Detection
To generate synthetic samples, we employ a quantized version of the Mistral-7B-instruct-v0.1 model [^2] https://huggingface.co/turboderp/Mistral-7B-instruct-exl2 to generate comments on a specific question $q$, using the following prompt:
A user in a discussion forum is debating other users about the following question: [q] The person is in favor about the topic in question. What would the person write? Write from the person’s first person perspective.
where "[q]" must be replaced with the question $q$. Similarly, to generate a negative sample, we replace "is in favor" with "is not in favor". As in the X-Stance dataset (Vamvas and Sennrich, 2020), we assign the two labels 0 and 1. We denote the question-specific synthetic dataset as:
@@
where half of the $M$ synthetic data samples have positive labels, i.e., are comments in favor for the posed question, while the other half is against. Since the dataset is in German, we translate the questions $q$ with a "NLLB-300M" (NLLB Team et al., 2022) translation model. The English answers from the Mistral-7B model are then translated back to German using the translation model.
We show the translation pipeline in Figure 9 along with translations of the questions chosen for our experiments in Table 2. We also attach the synthetic generated data in both English and German to the paper. In our experiments, this translation pipeline worked well even for complex topics in online political discussions, making our method applicable to any language where a good translation model exists. Overall, the generated dataset $\mathcal{D}_{\text{synth}}^{(q)}$ will be used in two ways: (i) to augment the existing dataset $\mathcal{D}^{(q)}$ in order to increase the amount of training data, and (ii) to detect the most informative samples in the unlabelled data pool, which is explained next.
3.3 Synthetic Data-driven Query By Comittee (SQBC)
In the typical active learning scenario, we have a set of unlabelled data, where the goal is to choose the most informative samples to label. Our approach is inspired by the two methods Query by Comittee (QBC) (Seung et al., 1992) and Contrastive Active Learning (CAL, Margatina et al. (2021)): instead of using QBC’s ensemble of experts and the KL-divergence based information score in CAL, we directly use the synthetic data and its labels to produce classification scores for the unlabelled data using $k$ nearest neighbors. The most informative samples are then the data points with the most indecisive scores. Since we query a synthetic dataset instead of an ensemble of experts (as in QBC), we call our method Synthetic data-driven QBC or short SQBC.
Our method performs three steps:
(1) Generate the embeddings.
Given some embedding function $g:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{E}$, we generate embeddings for the unlabelled data,
$E=\big{\{}e^{(i)}\big{\}}_{i=1}^{I}=\big{\{}g(q,x^{(i)})\big{\}}_{i=1}^{I}$ (2)
and for the labelled synthetic data
$E_{\text{synth}}=\big{\{}e_{\text{synth}}^{(m)}\big{\}}_{m=1}^{M}=\big{\{}g(q,x_{\text{synth}}^{(m)})\big{\}}_{m=1}^{M}.$ (3)
Note that $q$ is the question for which we generate the synthetic data and for which we want to detect the most informative samples. If obvious from the context, we often omit the superscript $(q)$.

### v2

@@ Abstract
Stance detection holds great potential to improve online political discussions through its deployment in discussion platforms for purposes such as content moderation, topic summarization or to facilitate more balanced discussions. Typically, transformer-based models are employed directly for stance detection, requiring vast amounts of data. However, the wide variety of debate topics in online political discussions makes data collection particularly challenging. LLMs have revived stance detection, but their online deployment in online political discussions faces challenges like inconsistent outputs, biases, and vulnerability to adversarial attacks. We show how LLM-generated synthetic data can improve stance detection for online political discussions by using reliable traditional stance detection models for online deployment, while leveraging the text generation capabilities of LLMs for synthetic data generation in a secure offline environment. To achieve this, (i) we generate synthetic data for specific debate questions by prompting a Mistral-7B model and show that fine-tuning with the generated synthetic data can substantially improve the performance of stance detection, while remaining interpretable and aligned with real world data. (ii) Using the synthetic data as a reference, we can improve performance even further by identifying the most informative samples in an unlabelled dataset, i.e., those samples which the stance detection model is most uncertain about and can benefit from the most. By fine-tuning with both synthetic data and the most informative samples, we surpass the performance of the baseline model that is fine-tuned on all true labels, while labelling considerably less data.
@@ 1 Introduction
With the recent advent of powerful generative Large Language Models (LLMs) such as ChatGPT, Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023), new ways of performing stance detection have opened up via zero-shot or chain-of-thought prompting. This is especially important in the area of online political discussion where topics are complex and labelled data is hard to come by. At the same time, an ever important use case in online political discussions is being able to use stance detection for an ongoing discussion to, e.g., suggest suitable comments for engagement between participants (Küçük & Can, 2020; Behrendt et al., 2024). In the case of LLMs, while strong at analysing complex topics and at open-ended text generation, explicit classification can be inconsistent (Cruickshank & Xian Ng, 2023), they are prone to biases (Ziems et al., 2023) and open to adversarial attacks (Greshake et al., 2023). More traditional stance detection models based on, e.g., BERT (Devlin et al., 2019) are task-specific and therefore consistent in their output, however they need large amounts of labelled data (Mehrafarin et al., 2022; Vamvas & Sennrich, 2020) to perform well.
@@
Figure 1: We investigate the use of LLM-generated synthetic data for stance detection in online political discussions. (A) We generate synthetic data  for specific questions using a Mistral-7B model. The synthetic data is then used to fine-tune the stance detection model. We show that fine-tuning with synthetic data improves the performance of the model, since the synthetic data is roughly faithful to the real data’s  underlying distribution. However, some real world samples  cannot be captured by the synthetic data. (B) We therefore use the synthetic data to identify the most informative samples  in the unlabelled real data pool, which are better off labelled by human experts. Combining the synthetic data with the manually labelled most informative samples improves the performance of the model even further.
In this work, we combine both traditional stance dectection and LLMs to get the best of both worlds. For stance dectection, we use BERT as a lightweight stance detection model that produces fast and consistent output given the data it has been fine-tuned on. To address the issue of needing large amounts of data, we propose to generate synthetic data with an LLM to augment the stance detection model for fine-tuning. This allows us to leverage LLMs in an offline setting to enhance classical stance detection models, which are better suited and safer for use in an online setting. Furthermore, we show that the synthetic data allows us to gain insights about the real world data. We show that the synthetic data generated by the LLM can serve as a reference distribution for stance detection, since the LLM is able to generate high quality samples which fall distinctly into the respective stance classes. This has two benefits: (i) we can analyse the data for potential biases by comparing the alignment of the synthetic data distribution to the real world data distribution. (ii) we can tackle the issue of extracting ambiguous samples that are difficult for the model to classify and therefore deteriorate its performance. Since the synthetic data partitions the space between both classes, ambiguous samples that the stance detection model cannot classify properly can be identified as lying in between the two classes. These samples can then be labelled manually. Fine-tuning with these additional samples, outperforms the model that is fine-tuned on all true labels alone, while only having labelled a subset of samples in the unlabelled data pool. We illustrate our method in Figure 1.
We view stance detection as a binary classification problem (favor or against), where we explore the following questions:
(Q1) Does fine-tuning a stance detection model with synthetic data improve stance detection performance? We first analyse whether fine-tuning the BERT model with synthetic data improves stance detection and show that this approach almost reaches the model trained with all true labels and is superior to using zero-shot Mistral-7B for the complex topics in online political discussions. We reveal that a stance detection model can be tailored to a certain topic with only synthetic data.
(Q2) How does the synthetic data improve performance and does it align with real world data? Our second question analyses the generated synthetic data. We analyse how well it aligns with the real world training data by visualising the T-SNE projected embeddings of the stance detection model and by comparing the entropy distribution of the synthetic data to the real data. We find that the synthetic data aligns well with the real data, indicating that the LLM is able to generate comments for both stances while introducing minimal further bias.
(Q3) Can we further improve the model by using the synthetic data as labelled reference distribution for active learning? The synthetic data allows us to identify unlabelled real data samples that improve the model even further through active learning. Due to the canonical nature of the synthetic data, we are able to extract real word samples for human labelling that are difficult (ambiguous) for the model to classify. We do this, by determining the $k$-nearest synthetic neighbours of the real data. The stance detection model is fine-tuned jointly with these samples and the synthetic data, where we surpass the baseline model even when it is fine-tuned on all true labels, while labelling considerably less data manually.
@@ 2 Background
Stance detection, a sub-task of sentiment analysis (Romberg & Escher, 2023) and opinion mining (ALDayel & Magdy, 2021), aims to automatically identify an author’s stance (favor, against, or neutral) towards a discussed issue or target. In online political discussions, this involves determining if the contribution in question is for or against a topic like tax increases. Stance detection has been identified as an important task for improving discussion summarization (Chowanda et al., 2017), detecting misinformation (Hardalov et al., 2022), and evaluating opinion distributions in online political discussion and participation processes (Romberg & Escher, 2023). Stance detection is also used in recommender systems and discussion platforms (Küçük & Can, 2020). Still, due to its dependency on context, stance detection is a highly challenging task. Identifying stance requires understanding both the question and the contributor’s position, complicated by users often deviating from the original question and discussing multiple topics in the same thread (Ziegele et al., 2014), leading to little usable training data. Some works in stance detection use graph convolutional networks to learn more out of the present data (Zhang et al., 2022; Li & Goldwasser, 2019). Recently, fine-tuning transformer-based models (Vaswani et al., 2017; Liu et al., 2022) to solve stance detection is a common practice, but training these models requires a large amount of annotated data, which for the large variety of questions in online political discussions is unfeasible to acquire.
@@
Active Learning. The aim of active learning is to minimize the effort of labelling data, while simultaneously maximizing the model’s performance. This is achieved by selecting a query strategy that chooses the most interesting samples from a set of unlabelled data points, which we refer to as most informative samples. These samples are then passed to, e.g., a human annotator for labelling. There exist many different query strategies such as Query By Comittee (QBC, (Seung et al., 1992)), Minimum Expected Entropy (MEE, Holub et al. (2008) or Contrastive Active Learning (CAL, Margatina et al. (2021)). By actively choosing samples and asking for the correct labelling, the model is able to learn from few labelled data points, which is advantageous especially when annotated datasets are not available. Within the domain of political text analysis, many different tasks lack large amounts of annotated data. It has been already shown in the past that these tasks can benefit from the active learning: e.g., stance detection (Kucher et al., 2017), topic modeling (Romberg & Escher, 2022), speech act classification (Schmidt et al., 2023) or toxic comment classification (Miller et al., 2020). In this work, we examine how LLM-generated synthetic data can be used instead of real labelled data to select the most informative samples to be manually labelled.
@@
Stance detection and synthetic data generation with LLMs. Recent work has shown that synthetic data generated from LLMs can be used to improve the performance of a model on downstream tasks. Møller et al. (2023) showed that synthetic data can be used to improve the performance of a model on downstream classification tasks by comparing the performance of a model finetuned on LLM-generated data to crowd annotated data. In many cases the model finetuned on LLM-generated data outperforms the model finetuned on crowd annotated data. Mahmoudi et al. (2024) study the use of synthetic data for data augmentation in stance-detection. The authors use GPT-3 to generate synthetic data for a specific topic with mixed results due to the inability of GPT-3 to generate good data samples. In our work,we use a newer LLM model, Mistral-7B, which generates better synthetic data samples and show that we can generate synthetic data that matches the real data distribution. Veselovsky et al. (2023) analyse in which ways synthetic data is best generated for tasks like sarcasm detection and sentiment analysis. The authors reach the conclusion that grounding the prompts to generate the synthetic data to real samples helps improve the quality of the synthetic data. Similarly, Li et al. (2023) argue that subjectivity in the classification task determines whether synthetic data can be used effectively. It has been shown that LLMs can be used directly for stance detection such as (Cruickshank & Xian Ng, 2023), (Burnham, 2023) (Ziems et al., 2023). However, the general conclusion of these studies is that while LLMs are competitive with other transformer models such as BERT, especially for edge cases, they exhibit replication issues. (Burnham, 2023) also discuss the posibility of pre-training models on more specific data to improve the generalisation capability of the model. Ziems et al. (2023) highlight the potential biases that can emerged in open ended generation tasks and classification performance varies depending on how representative the training data is. We therefore focus on using LLMs to generate synthetic data to solve key challenges in stance detection such as the lack of available data for specific topics and labelling large amounts of data, rather than using LLMs directly for the task.
@@ 3 Method
In line with our declared contributions, we present our core ideas to improve the performance of the stance detection model: (i) To fine-tune the model with synthetic data, we first define the synthetic dataset and show our pipeline to generate it. The baseline model is then further fine-tuned on the synthetic data. (ii) We then present our synthetic data-based approach to identify the most informative samples from a batch of unlabelled data, where we propose a synthetic extension to the QBC (Query by Commitee, (Seung et al., 1992)) method, where the synthetic data act as an ensemble of experts. As described in Section 1 and Figure 1, we use the synthetic data as reference distribution to identify the most informative samples in the unlabelled data pool. The idea is that for ambigouous samples the k-synthetic nearerst neighbours are split in their labels and therefore lie on the decision boundary of the model.
@@ 3.1 Preliminaries
Political discussions are typically centered around questions $q\in\mathcal{Q}$ (sometimes also called issues or targets). For stance detection, we usually have for each of these questions $q$ a set of labelled data $\mathcal{D}^{(q)}=\{(x^{(i)},y^{(i)})\}_{i=1}^{I}$ where $x^{(i)}\in\mathcal{X}$ is a statement (or comment) and $y^{(i)}$ is the stance of the statement, with $y^{(i)}\in\{0,1\}=\mathcal{Y}$. Note, that we use the notation $\mathcal{D}^{(q)}$ for labelled and for unlabelled datasets (then the labels are ignored). We view the stance detection model as a binary classification function $f:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{Y}$, where we included the question as input to provide context. The stance detection model such as BERT (Devlin et al., 2019) is fine-tuned by minimizing the cross-entropy loss between the predicted labels $\hat{y}^{(i)}=f(q,x^{(i)})$ and the actual labels $y^{(i)}$.
@@ 3.2 Generating Synthetic Data for Stance Detection
To generate synthetic samples, we employ a quantized version of the Mistral-7B-instruct-v0.1 model to generate comments on a specific question $q$, using the following prompt: A user in a discussion forum is debating other users about the following question: [q] The person is in favor about the topic in question. What would the person write? Write from the person’s first person perspective.
where "[q]" must be replaced with the question $q$. Similarly, to generate a negative sample, we replace "is in favor" with "is not in favor". As in the X-Stance dataset (Vamvas & Sennrich, 2020), we assign the two labels 0 and 1. We denote the question-specific synthetic dataset as:
@@
where half of the $M$ synthetic data samples have positive labels, i.e., are comments in favor for the posed question, while the other half is against. Since the dataset is in German, we translate the questions $q$ with a "NLLB-300M" (NLLB Team et al., 2022) translation model. The English answers from the Mistral-7B model are then translated back to German using the translation model. We also tried other similar sized open source LLMs (Llama, Openassistant, Falcon), but found that only the Mistral-7B model produced sensible comments.
Overall, the generated dataset $\mathcal{D}_{\text{synth}}^{(q)}$ will be used in two ways: (i) to augment the existing dataset $\mathcal{D}^{(q)}$ in order to increase the amount of training data, and (ii) to detect the most informative samples in the unlabelled data pool, which is explained next.
3.3 Getting the Most Informative Samples: Synthetic Query By Comittee
To identify the ambiguous (most informative) samples as described in (Q3) we take from two active learning methods: Query by Comittee (QBC) (Seung et al., 1992) and Contrastive Active Learning (CAL, Margatina et al. (2021)). Instead of using QBC’s ensemble of experts and the KL-divergence based information score in CAL, we directly use the synthetic data and its labels to identify ambigous samples using $k$ nearest neighbors. The most informative samples are then the data points with the most indecisive scores. Synthetic Query by Comitte (SQBC) consists of three steps:
(1) Generate the embeddings. Given some embedding function $g:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{E}$, we generate embeddings for the unlabelled data, $E=\big{\{}e^{(i)}\big{\}}_{i=1}^{I}=\big{\{}g(q,x^{(i)})\big{\}}_{i=1}^{I}$ and for the labelled synthetic data $E_{\text{synth}}=\big{\{}e_{\text{synth}}^{(m)}\big{\}}_{m=1}^{M}=\big{\{}g(q,x_{\text{synth}}^{(m)})\big{\}}_{m=1}^{M}.$ Note that $q$ is the question for which we generate the synthetic data and for which we want to detect the most informative samples. If obvious from the context, we often omit the superscript $(q)$.